Unemployment among high school graduates is quite high due to a recession. City Community College is considering a new program to help young people get the training they need to be more employable. The college has collaborated with the municipal hospital to build a nurse’s aide program, a 1-year program that would lead to immediate employment. Initial financial analysis indicates that the total fixed cost of the program will be $200,000, which includes the cost of a 1-year rental of the facilities plus utilities, insurance, and administrative costs. The variable cost and step cost together are $10,000 per student, which pays for faculty salary, student lunches, and teaching materials and textbooks. The state kicks in $5,000 per student, and the tuition based on market analysis is $6,000. Given that the nurse’s aide program has never been offered in the region before and will be under financial pressure due to current funding cuts, the college’s board of trustees would like to know how many students would need to be enrolled for the program to break even. The board also wants to know what other options it will have to mitigate these financial issues if the expected enrollment is below the break even point, given the high profile of the program at a time when employment and economic recovery are critical.Consider two possibilities: (1) Funding is limited to $100,000, or (2) with the benefit of efficiencies, variable and step costs can be reduced to $9,500. You are required to do a financial analysis of the proposed program. Please provide a spreadsheet solution and a written explanation of your approach.
Answer #1

The Break Even Point is a point where total revenue is equal to total expenses

Revenue = Fixed Cost + Variable Cost

P*n= F+V*n

P= Govt Funding + Tuition fees

where P is price per unit, n is number of unit, V is variable cost per unit, F is Total fixed cost

n= F/(P-V)

The Break-even for given scenario is calculated as:

n = 200,000 ( (5,000+6,000)-10,000) = 200

Thus for break even, enrollment of 200 students is required

Possibility 1)

Funding is limited to 100,000

Revenue = Fixed Cost + Variable Cost

P*n= F+V*n

Now P= G+T

where G is govt contribution per student and T is tuition fees per student

Govt Funding + Tuition fees Collection = F+V*n

100,000 + 6,000*n = 200,000 + 10,000*n

6,000*n - 10,000*n = 200,000 - 100,000

Thus it is not possible to sustain with current tuition fees of 60,000.

It must be atleast 10,001 to be able to sustain with enrollment being required will be 100,000

Possibility 2)

Variable cost reduced to 9,500 per student

Govt Funding + Tuition fees Collection = F+V*n

5,000*n + 6,000*n = 200,000 + 9,500*n

11,000*n = 200,000 + 9,500*n

1,500*n = 200,000

n= 133.33 = 134

Thus if Variable cost is reduced to 9,500 per student then enrollment required to break-even will reduce from initial 200 to 134